Pauls age is 8 years less than 5 times sarahs age. the sum of their ages is 64. what is pauls age??? heeellpppp

Answer by stanbon(75887)

Pauls age is 8 years less than 5 times sarahs age.
the sum of their ages is 64. what is pauls age???
------
Equations:
p = 5s-8
p + s = 64
-------
Rearrange:
p - 5s = -8
5p+ 5s = 5*64
---------
Add and solve for "p"::
6p = 312
p = 52 years (paul's age now)
------------------------------------
Cheers,
Stan H.
----------------